OpenStudy (cheska_p):

First derivative cot(x)cos(x). The correct answer is -cosx-cotxcscx, but I don't know how.

Nnesha (nnesha):

how would you start ??

Nnesha (nnesha):

\[\large\rm \cot(x) \cdot \cos(x)\] there are two functions so apply the product rule ?\[\large\rm y=f(x) \cdot g(x) ~~~~~~~~~y'= f'(x) \cdot g(x) + f(x) \cdot g'(x)\]

OpenStudy (cheska_p):

Yes, when I use that method I get -csc^2(x)cos(x)-cot(x)sin(x)

Nnesha (nnesha):

yep looks good now rewrite cot and csc^2 in terms of cos ., sin :)

Nnesha (nnesha):

\[\large\rm -\color{reD}{\csc^2(x)}\cos(x)-\color{blue}{\cot(x)}\sin(x)\] \[\large\rm -\color{red}{\frac{ 1 }{ \sin^2x }} \cos(x) - \color{blue}{\frac{\cos x}{\sin x}} \cdot \sin(x)\] sin^2x is same as (sinx times sin x) so \[\large\rm -\color{red}{\frac{ 1 }{ sinx \cdot \sin x }} \cos(x) - \color{blue}{\frac{\cos x}{\sin x}} \cdot \sin(x)\] now simpify

Nnesha (nnesha):

let me know if you have question

OpenStudy (cheska_p):

I'm not entirely sure how to ask this question, but I'll try. Why can the sin(x) on right side of the subtraction sign simplify the sin(x) on the left side of the subtraction sign? I thought it couldn't do that.

OpenStudy (cheska_p):

@Nnesha

Nnesha (nnesha):

http://prnt.sc/aingsr are you talking about these sin ratios ?

OpenStudy (cheska_p):

@Nnesha yes I am

Nnesha (nnesha):

ohh yea we can't do that bec 1/sin^2 is a different term just like in this example we can't just cancel out y . first we need to find common denominator \[\large\rm \frac{1}{y}- y\]

Nnesha (nnesha):

\[\large\rm -\color{red}{\frac{ 1 }{ sinx \cdot \sin x }} \cos(x) - \color{blue}{\frac{\cos x}{\cancel{\sin x}}} \cdot \cancel{\sin(x)}\] you can cancel out sinx in 2nd term bec its a multiplication i don't know if that make sense .. hm

OpenStudy (cheska_p):

I just realized something, the way you're solving the problem doesn't give me the correct answer either

OpenStudy (cheska_p):

@Nnesha I figured out, you weren't wrong I was just visualizing my answer incorrectly. Thank you for your help!

Nnesha (nnesha):

aw glad to hear that if i split this fraction into two fractions then it would be easy to understand \[\large\rm -\color{red}{\frac{ 1 }{ sinx }\cdot \frac{1}{\sin x }} \cdot \cos(x) - \color{blue}{\frac{\cos x}{\cancel{\sin x}}} \cdot \cancel{\sin(x)}\] \[\large\rm -\color{red}{\frac{ 1 }{ sinx }\cdot \frac{\color{black}{cos x }}{\sin x }} - \color{blue}{\frac{\cos x}{\cancel{\sin x}}} \cdot \cancel{\sin(x)}\] :)

Nnesha (nnesha):

& my pleasure :)